# Learning in a Multifield Coherent Ising Machine

### v1

@@ Introduction
The interaction springs between computational degrees of freedom $\vec{\psi}_{i}$ and $\vec{\psi}_{j}$ are determined by the long term memory (weights) of the model, encoded in the spin texture of the $\sigma$ field. In our model, this dependence arises through a nonlinear interaction $H_{I}$ between each pair of nearest-neighbor sites $i$ and $j$, with the form:
$H_{I}=\lambda(t)\left[c_{0}+(\sigma_{i}-\sigma_{j})^{2}\right]\vec{\psi}_{i}^{T}\vec{\psi}_{j},$ (5)
This term induces an effective coupling spring $c_{eff}$ between modes $\psi_{x,i}$ and $\psi_{x,j}$, as well as between modes $\psi_{y,i}$ and $\psi_{y,j}$; with values $c_{eff}=\lambda(t)c_{0}$ when the spins $\sigma_{i}$ and $\sigma_{j}$ are aligned, and $c_{eff}=\lambda(t)(c_{0}+2A_{\sigma}^{2})$ when $\sigma_{i}$ and $\sigma_{j}$ are anti-aligned (See Fig. 1 for a visual representation of this effect). Here, $A_{\sigma}$ is the amplitude of oscillation of the Ising mode. During inference, the coupling strength $\lambda(t)$ is set to a fixed value $\lambda_{0}$; while at each training iteration it is momentarily set to 00 so that sites exactly follow the single-site learning dynamics described earlier. The term $c_{0}$ ensures that there is some coupling even when the spins $\sigma_{i}$ and $\sigma_{j}$ are aligned—thus preventing a large number of spin configurations from collapsing to zero transmissivity. Although this work corresponds to a tight-binding model, in an experimental setting, the potential could be realized by a set of appropriately placed cubic springs (see Appendix C).
Figure 4: Iris flower classification using a multifield coherent Ising machine. (a) The Iris dataset consists of the petal length $l_{p}$, petal width $w_{p}$, sepal length $l_{s}$ and sepal width $w_{s}$ for 150 flowers belonging to three classes (iris setosa, iris virginica, iris versicolor). (b) The features are injected into the multifield coherent Ising machine, by encoding them in the amplitude of harmonic excitations at the computational frequency $\omega_{c}$. The output is taken at the central site. Positive and negative copies of the signals are applied, as the lattice cannot perform subtractions. The full model consists of three multifield Coherent Ising Machines, corresponding to each of the model classes. The machines learn to produce a high amplitude when excited by a sample of their corresponding class. (c) Evolution of the mean classification accuracy during training. The shaded area corresponds to one standard deviation. (d) Histograms of the classification accuracy computed on an untrained lattice (blue), and after 20 (orange) and 98 (green) training iterations. The training times corresponding to the histograms are indicated as solid dots in panel (c).
Due to the interaction term $H_{I}$, during inference ($\lambda(t)=\lambda_{0}$ and $\mu(t)=0$), the dynamics of the $\vec{\psi}$ field can be understood as a spin-dependent effective spring network, whose steady-state amplitude response is given by the linear system $K(s)A_{\psi_{x/y}}=f_{\psi_{x/y}}$. Here, $K(s)$ is the spin-dependent stiffness matrix, whose diagonal components are $K_{i,i}=i\omega_{c}^{2}/Q_{c}$ and the off-diagonal components are given by $K_{i,j}=-c_{eff}(s_{i},s_{j})$ when $i$ and $j$ are nearest neighbors, and zero otherwise. We choose $\lambda_{0}$ and $c_{0}$ so that the coupling stiffness is $10$ times the damping rate $\omega^{2}_{c}/Q$ for anti-aligned spins, and $2.5$ times the damping rate for the aligned case. Here, we have introduced the variable $s_{i}\in\{+1,-1\}$ to label the two possible phase oscillation states of $\sigma_{i}$. Gauge freedom means that the specific labeling is irrelevant as long as it is consistent between sites. We use this linearized model to study the expressivity of a square lattice with $44x$ 4 sites (Fig. 3a)—which output amplitude responses $A_{o}$ can be encoded as a spin configuration $s$. We do so by computing the transmitted amplitude associated with every spin configuration and calculating an effective density of states (Fig. 3b,c). We observe that the lattice can encode a range of mostly positive transmissivities, with two small gap regions (Fig. 3b).
Our lattice model can learn to approximate a given transmission amplitude (Fig. 3d, e), as long as the target amplitude can be expressed by a spin configuration. The learning protocol is as follows: We apply a harmonic force with magnitude $F_{i}$ and frequency $\omega_{i}$ at both clamped $\psi_{x}$ and free $\psi_{y}$ degrees of freedom of the the top-left site, while prescribing the displacement of the clamped mode $\psi_{y}$ at bottom-right site to a target value $A_{T}$. During this procedure, we keep $\lambda(t)=\lambda_{0}$ and $\mu(t)=0$, waiting for a sufficient time ($\Delta T\gg Q_{c}/\omega_{c}$) so that the $\vec{\psi}$ field equilibrates. Then, we set $\lambda(t)=0$ and apply a Gaussian learning pulse $\mu(t)$ to every site—during this protocol, the contrast $\psi_{x-y}$ remains large owing to the long lifetime of the computational modes. For every target output amplitude $A_{T}$, we repeat this procedure $n$ times, starting from a random spin configuration.
@@
We illustrate the capabilities of the proposed system on the Iris classification test [30]. The dataset consists of measurements of the petal length $l_{p}$, petal width $w_{p}$, sepal length $l_{s}$ and sepal width $s_{w}$ (Fig. 4a) for a set of 150 flowers belonging to the species iris setosa, iris virginica, and iris versicolor. The classifier consists of a system of three 13x13 lattices with one-hot encoding; the lattice with the highest output amplitude, measured at the central site, will be the inferred flower species. We apply the geometric features ($l_{p}$, $w_{p}$, $l_{s}$, $w_{s}$) by encoding them in the amplitudes of a set harmonic excitation forces at the frequency of the computational modes. These signals are applied at sites through the boundary of the lattice (Fig. 4b), acting on both free $\psi_{x}$ and clamped $\psi_{y}$ computational modes. Since the lattices cannot produce large negative transmissivities (Fig. 3b), we apply both positive and negative copies of each feature signal.
The training starts from three lattices with random spin configurations. Then, we apply the forces encoding the features corresponding to a randomly chosen flower to each of the three lattices. We set $\lambda=\lambda_{0}$ and $\mu=0$, and clamp the output site $\psi_{y,center}$ to $55$ for the lattice corresponding to the current type of flower, and to $-35$ for the lattices corresponding to the other flowers. Then, we apply a learning protocol to all lattices, consisting of setting $\lambda=0$ and $\mu(t)$ to a Gaussian pulse, inducing bit flips when clamped and free sectors disagree. We simulate the training protocol by computing the steady-state amplitudes using the linearized model, and then flipping the bits with probabilities given by Eq. 4 with $\alpha=10$ and $\psi_{T}=1.5$. We repeat this training for 2000 iterations, observing that the system starts from an accuracy of $33.2\pm 22.1\,\%$ , as one would expect from random choice. As more training iterations are performed—each consisting of a single, randomly-selected flower sample, the accuracy increases to $71.9\pm 9.8\,\%$, more than halving the error rate, and significantly outperforming random choice. While this accuracy is lower than prior works on physical learning, we use a much simpler training method and network architecture. We expect that complex hierarchical networks trained with complex protocols should be able to reach performances comparable with the state-of-the-art.
@@ Appendix A Determination of the parametric resonance regions
This is the expression used to determine the region of self-resonance in Fig. 2.
@@ Appendix B Numerical methods
B.1 Simulation of the learning process
We simulate this process by computing the equilibrium amplitudes of $\vec{\psi}$ using the linearized model, and then determining the flip probability using the single-site flip probability (Eq. 4)—which provides an exact result as we switch off the inter-site potential during coupling.
@@ 2 $(0,0,1)$ $(0,0,1)$
Of the seven interaction potential terms in Table 1, the first two introduce the spin-dependent inter-site interactions between modes $\psi_{x}-\psi_{x}^{\prime}$ and $\psi_{y}-\psi_{y}^{\prime}$ respectively; the next four terms compensate for spin-dependent changes in the natural frequency of the computational degrees of freedom, while the last term compensates for spin-dependent changes in the Ising degrees of freedom. These extra terms are requierd to cancel the contribution of cross-products in Eq. 12, that do not exist in Eq. C. Only the first two terms determine the flow of information, as they couple computational degrees of freedom between sites. In absence of the compensating terms, the system may still be capable of learning. However, in this case, the output phase becomes dependent on the spin texture, which complicates the learning process.

### v2

@@ Results
In machine learning models, the role of the weights is to determine what computation is carried out, by setting the strength of the coupling between neurons. Consequently, in a self-learning physical system, storing weight and computational degrees of freedom separately is not sufficient; we must also ensure that the weights influence the computation that is being performed. In our metamaterial, this influence is established through a nonlinear interaction $H_{I}$ acting between each pair of degrees of nearest-neighbor sites $i$, $j$. It is governed by an energy of the form
$H_{I}=-\lambda(t)\left[c_{0}+(\sigma_{i}-\sigma_{j})^{2}\right]\vec{\psi}_{i}^{T}\vec{\psi}_{j},$ (6)
This term generates the effective coupling springs that connect neighboring computational degrees of freedom ($\psi_{x,i}\leftrightarrow\psi_{x,j}$ and $\psi_{y,i}\leftrightarrow\psi_{y,j}$). Depending on the phase of the weight degrees of freedom, these coupling springs can take two values (Fig. 1)—determined by $\lambda(t)=\lambda_{m}$ and $c_{0}$; if the weight field has the same phase of oscillation in both sites, the spring constant will be $k_{low}=\lambda_{m}c_{0}$, while if weights oscillate at opposing phases, the spring constant will be $k_{high}=\lambda_{m}(c_{0}+4<\sigma^{2}>)$. This effective spring model is valid as long as the computational amplitude remains small and the potential in Eq. 6 does not alter the weight amplitude.
Although each individual weight $\sigma_{i}$ is binary, in metamaterials composed of multiple sites (Fig. 3a) the transmitted energy between input and output can take many different values, depending on the global weight configuration. This can be seen in the density of states, that quantifies the number of weight configurations that result in a lattice transmissivity around a specific value (Fig. 3b,c). Consequently, such lattice can be used to approximate continuous transmissivities, as long as they fall within the range of expressible values.
To perform inference with the metamaterial, we encode the input in the amplitude of a harmonic force, applied to both computational degrees of freedom, $\psi_{x,input}$ and $\psi_{y,input}$, at the input site. We let the lattice reach a steady-state, and read the result in the amplitude of the output computational degree of freedom, $\psi_{x,output}$. During this process, we keep $\lambda(t)$ at its inference value $\lambda_{m}$ and $\mu(t)$ at zero. To perform a training iteration, in addition to the input forces, we also clamp the output site $\psi_{y,output}$ to the desired value. Then, once the lattice reaches a steady-state, we perform a weight update by increasing $\mu(t)$ according to the protocol discussed in the previous section. During the weight update, we momentarily set $\lambda(t)=0$, turning the lattice into a set of disconnected sites. Under these conditions, the learning protocol $\mu(t)$ introduces a probabilistic weight flip, whose initial condition is given by the steady-state amplitude of the computational degrees of freedom. Figure 3d,e illustrates the training of a metamaterial to match a specific signal transmissivity. At every training iteration, the expected value of the output energy gets closer to the target value, while the variance between different runs decreases.
@@
Figure 4: Iris flower classification using a multifield coherent Ising machine. (a) Flower features contained in the Iris dataset. (b) The features are injected into the multifield coherent Ising machine, by encoding them in the amplitude of harmonic excitations at the computational frequency $\omega_{c}$. The output is taken at the central site. Positive and negative copies of the signals are applied, as the lattice cannot perform subtractions. The full model consists of three multifield Coherent Ising Machines, corresponding to each of the model classes. The machines learn to produce a high amplitude when excited by a sample of their corresponding class. (c) Evolution of the mean classification accuracy during training. The shaded area corresponds to one standard deviation. (d) Histograms of the classification accuracy computed on an untrained lattice (blue), and after 20 (orange) and 98 (green) training iterations. The training times corresponding to the histograms are indicated as solid dots in panel (c).
@@
We illustrate the capabilities of the proposed system on the Iris classification test [33]. The dataset contains measurements of the petal length $l_{p}$, petal width $w_{p}$, sepal length $l_{s}$ and sepal width $s_{w}$ (Fig. 4a) for a set of 150 flowers belonging to the species iris setosa, iris virginica, and iris versicolor. The classifier consists of a system of three 13x13 lattices; the lattice with the highest output amplitude, measured at the central site, will be the inferred flower species. We encode the geometric features ($l_{p}$, $w_{p}$, $l_{s}$, $w_{s}$) in the amplitudes of a set harmonic excitation forces acting on each lattice. These signals are applied at sites through the boundary of the lattice (Fig. 4b), acting on both free $\psi_{x}$ and clamped $\psi_{y}$ computational modes. Since the lattices cannot produce large negative transmissivities (Fig. 3b), we apply both positive and negative copies of each feature signal.
The training process starts by initializing each of the three lattices with a random weight configuration. Then, for every learning iteration, we subject all lattices to the forces encoding the features of a randomly chosen flower. We clamp the output site $\psi_{y,center}$ to $55$ for the lattice corresponding to the current type of flower, and to 35-35 for the lattices corresponding to the other flowers. Finally, we apply a weight update protocol to all lattices, consisting of setting the coupling $\lambda$ to zero, and changing $\mu(t)$ following a Gaussian pulse. This process is repeated for 2000 iterations.
We observe that the system starts from an accuracy of $33.2\pm 22.1\,\%$ , as one would expect from random choice. As more training iterations are performed, the accuracy increases to $71.9\pm 9.8\,\%$, providing a clear signature of physical learning. These results are remarkable as the system stores only 156 bits of information, contains only Kerr nonlinearities, and is directly trained by exposing it to the desired output. Using a more advanced training algorithm, where the learning process is stopped when the precision exceeds a threshold, we achieve $80.0\pm 7.0\,\%$ accuracy, with the best-performing runs reaching $87.3\,\%$. We hypothesize that larger, architected networks trained under more complex protocols, such those where the target value for every sample is adjusted according to the lattice output [23], the system could be able to reach performances comparable with electronic linear classifiers.
Discussion
We have shown that a metamaterial composed of a network of parametrically-driven, multimodal resonators can learn from examples—autonomously acquiring the capability to solve a classification task. The metamaterial training is based on a contrastive learning rule, which involves comparing two copies of the system—one left free, and one clamped to the desired output—and updating the model weights according to the difference between copies. However, unlike traditional contrastive learning, our model weights are binary and cannot be updated continuously. We circumvent this limitation by reformulating the learning rule as a probabilistic process, where the probability of a weight flip is a function of the mismatch between the two copies. This mismatch-dependent flip rate is reminiscent of evolutionary mechanisms in bacteria, where the mutation rate increases when the organism is under stress [34, 35, 36]. Similarly, our system increases the weight mutation rate at sites where its clamped and free copies disagree.
In our model, learning emerges from the interplay of symmetry, multistability, and noise, features not typically associated with computation. This work therefore contributes to the growing evidence that such physical effects hold significant, untapped potential for information processing [37, 38, 39, 40]. In fact, symmetry plays an essential role in the learning process: Clamping the $\psi_{y}$ breaks the symmetry between the two components of the computational field $\vec{\psi}$. The weight field, $\sigma$, then evolves to restore this broken symmetry over time. As symmetry is gradually restored, the free part of the system learns to mimic the clamped part, not by accessing the training data directly, but through persistent changes in the weight variable $\sigma$. This perspective suggests a novel symmetry-based design principle for self-learning materials, much like symmetry considerations guide the design of emergent condensed matter phenomena such as topological edge states [41]. The presence of multiple time scales is also an essential component. These are important in the brain [42] and can be replicated in artificial systems through parametric drive.
By providing a tractable physical model for emergent learning—analogous to the tight-binding Hamiltonians that describe topological insulators and superconductors—our work invites condensed matter-inspired questions about learning. For instance, one can now investigate the influence of disorder [43, 44], aging [45], and phase transitions [46, 47, 48] on learning behavior. As the constituent oscillators can be realized with components that operate at optical frequencies [49], consume little energy [50], or are integrated with very high densities [51], our model may inspire new directions to train machine learning models with physical computers.
@@ Appendix A Determination of the parametric resonance regions
This is the expression used to determine the region of self-resonance in Fig. 2. It should be noted that its value determines the minimum pumping that leads to self-resonance. Numerical simulations are conducted at twice this value, to ensure sufficient stability of the self-resonant state.
@@ Appendix B Numerical methods
Simulation of the learning process
[…]
For the learning examples of Fig. 3 and Fig. 4, we use the parameters $k_{high}=10\omega_{c}^{2}/Q_{c}$, $4k_{low}=k_{high}$, $\beta=0.273$ and $\psi_{T}=55$. In Fig. 3, we use the flip probability parameters $\beta=750$ and $\psi_{T}=0.16$, while in Fig. 4, the parameters are $\beta=0.273$ and $\psi_{T}=55$.
@@ 2 $(0,0,1)$ $(0,0,1)$
Of the seven interaction potential terms in Table 1, the first two introduce the spin-dependent inter-site interactions between modes $\psi_{x}-\psi_{x}^{\prime}$ and $\psi_{y}-\psi_{y}^{\prime}$ respectively; the next four terms compensate for spin-dependent changes in the natural frequency of the computational degrees of freedom, while the last term compensates for spin-dependent changes in the Ising degrees of freedom. These extra terms are required to cancel the contribution of cross-products in Eq. 13, that do not exist in Eq. 6. Only the first two terms determine the flow of information, as they couple computational degrees of freedom between sites. In absence of the compensating terms, the system may still be capable of learning. However, in this case, the output phase becomes dependent on the spin texture, which complicates the learning process.